# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ 2.5. Replication Package
To enable repeating our study, as well as verifying our findings, we publicly released a replication package on Zenodo (Danushi et al., 2024). The replication packages includes the intermediary artifacts produced while searching and selecting the studies to be considered, as well as their final classification. The selected studies can instead be accessed by relying on the bibliographic information provided by the references available in Table 2.
@@ 3. Overviewing the State-of-the-Art Solutions
(Ahmed, 2020) | ✓ | ✓ | sustainability guidelines | GP | WB | 2 | ✓ | ✓ | cloud computing | ✓ | ✓ | JA | 2020 | ||||||||||||
(Alofi et al., 2023) | ✓ | ✓ | sw improvement techniques | SB | BB | 4 | ✓ | ✓ | blockchain | ✓ | JA | 2023 | |||||||||||||
(Alofi et al., 2022) | ✓ | ✓ | sw improvement techniques | SB | BB | 4 | ✓ | ✓ | blockchain | ✓ | JA | 2022 | |||||||||||||
(Atkinson et al., 2014) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 3 | ✓ | cloud computing | ✓ | ✓ | JA | 2014 | |||||||||||
(Becker et al., 2016) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | JA | 2016 | |||||||||
(Beghoura et al., 2014) | ✓ | sw improvement techniques | ML | BB | 4 | ✓ | mobile app. | ✓ | CP | 2014 | |||||||||||||||
(Brownlee et al., 2021) | ✓ | sustainability guidelines | ER | BB | 4 | ✓ | ML | ✓ | ✓ | CP | 2021 | ||||||||||||||
(Chauhan and Saxena, 2013) | ✓ | ✓ | ✓ | sustainability guidelines | GP | BB | 1 | ✓ | ✓ | cloud computing | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2013 | ||||||||
(Condori-Fernández and Lago, 2018) | ✓ | sustainability guidelines | GP | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | CP | 2018 | |||||||||||||
(Dick et al., 2013) | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | ||||||||
(Dick et al., 2010) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | BC | 2010 | |||||||
(Fakhar et al., 2012) | ✓ | sw improvement techniques | SA | WB | 3 | ✓ | ✓ | cloud computing | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2012 | ||||||||||
(Fu et al., 2021) | ✓ | sustainability guidelines | GP | BB | 3 | ✓ | ✓ | comp. vis. | ✓ | CP | 2021 | ||||||||||||||
(Goaër, 2021) | ✓ | sw improvement techniques | SA | WB | 3 | ✓ | mobile app. | ✓ | ✓ | CP | 2021 | ||||||||||||||
(Goaër, 2023) | ✓ | sw improvement techniques | SA | WB | 3 | ✓ | ✓ | mobile app. | ✓ | CP | 2023 | ||||||||||||||
(Goaër and Hertout, 2022) | ✓ | sw improvement techniques | SA | WB | 4 | ✓ | ✓ | mobile app. | ✓ | CP | 2023 | ||||||||||||||
(Heithoff et al., 2023) | ✓ | ✓ | sw improvement techniques | MD | BB | 1 | ✓ | ✓ | cyb. phys. sys. | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2023 | |||||||||
(Hoesch-Klohe et al., 2010) | ✓ | ✓ | ✓ | sw improvement techniques | MD | BB | 4 | ✓ | generic sw | ✓ | CP | 2010 | |||||||||||||
(Ibrahim et al., 2021) | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | JA | 2021 | |||||||||||||
(Ibrahim et al., 2022) | ✓ | ✓ | reference models | RM | BB | 3 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2022 | |||||||||
(Jayanthi et al., 2021) | ✓ | ✓ | ✓ | reference models | RM | BB | 1 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2021 | ||||||||
(Jebraeil et al., 2017) | ✓ | sw improvement techniques | MD | BB | 4 | ✓ | ✓ | cloud computing | ✓ | CP | 2017 | ||||||||||||||
(Johann et al., 2011) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2011 | |||||||
(Karunakaran and Rao, 2013) | ✓ | sw improvement techniques | MD | BB | 4 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | |||||||||||
(Kern et al., 2013) | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | generic sw | ✓ | CP | 2013 | |||||||||||||
(Kern et al., 2015) | ✓ | ✓ | measurement solutions | ET | BB | 4 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | BC | 2015 | ||||||||||
(Kern et al., 2018) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 3 | ✓ | generic sw | ✓ | JA | 2018 | ||||||||||||
(Kharchenko and Gorbenko, 2013) | ✓ | sustainability guidelines | GP | BB | 2 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | |||||||||||
(Kipp et al., 2011) | ✓ | measurement solutions | ET | BB | 2 | ✓ | HPC | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2011 | |||||||||||
(Kipp et al., 2012) | ✓ | measurement solutions | ET | BB | 3 | ✓ | HPC | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2012 | |||||||||||
(Koçak and Alptekin, 2019) | ✓ | ✓ | sw improvement techniques | DM | BB | 2 | ✓ | generic sw | ✓ | ✓ | CP | 2013 | |||||||||||||
(Lago, 2019) | ✓ | sw improvement techniques | DM | BB | 3 | ✓ | ✓ | generic sw | ✓ | CP | 2019 | ||||||||||||||
(Lago and Jansen, 2010) | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | CP | 2011 | |||||||||||
(Lami and Buglione, 2012) | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2012 | ||||||||||
(Lami et al., 2014) | ✓ | reference models | RM | BB | 2 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2014 | |||||||||||
(Lami et al., 2013) | ✓ | reference models | RM | BB | 2 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | |||||||||||
(Lasla et al., 2022) | ✓ | sw improvement techniques | DP | BB | 4 | ✓ | blockchain | ✓ | JA | 2022 | |||||||||||||||
(Mahmoud and Ahmad, 2013) | ✓ | reference models | RM | WB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2013 | ||||||||||
(Mancebo et al., 2021) | ✓ | ✓ | ✓ | sw improvement techniques | TE | WB | 4 | ✓ | generic sw | ✓ | BC | 2021 | |||||||||||||
(Mehra et al., 2023) | ✓ | sustainability guidelines | ER | WB | 5 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | CP | 2023 | ||||||||||||
(Mehra et al., 2022) | ✓ | reference models | RM | BB | 3 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2015 | |||||||||||
(Mohankumar and Kumar, 2016) | ✓ | measurement solutions | AT | BB | 2 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2022 | |||||||||||
(Moshnyaga, 2013a) | ✓ | reference models | RM | BB | 2 | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2016 | |||||||||||
(Moshnyaga, 2013b) | ✓ | measurement solutions | ET | BB | 4 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | ||||||||||
(Moshnyaga, 2013c) | ✓ | measurement solutions | ET | BB | 4 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | ||||||||||
(Naumann et al., 2011) | ✓ | ✓ | measurement solutions | ET | BB | 4 | ✓ | ✓ | mobile app. | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2013 | |||||||||
(Naumann et al., 2015) | ✓ | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2001 | |||||||
(Noureddine et al., 2012) | ✓ | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | BC | 2015 | ||||||||
(Pa et al., 2017) | ✓ | ✓ | measurement solutions | ET | BB | 4 | ✓ | generic sw | ✓ | ✓ | CP | 2012 | |||||||||||||
(Pan and Rajan, 2022) | ✓ | measurement solutions | AT | BB | 3 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2017 | ||||||||||
(Patterson et al., 2022) | ✓ | sustainability guidelines | GP | BB | 4 | ✓ | ML | ✓ | CP | 2022 | |||||||||||||||
(Penzenstadler and Femmer, 2013) | ✓ | sustainability guidelines | GP | BB | 4 | ✓ | ✓ | ML | ✓ | JA | 2022 | ||||||||||||||
(Ponsard et al., 2018) | ✓ | ✓ | reference models | RM | BB | 3 | ✓ | ✓ | generic sw | ✓ | CP | 2013 | |||||||||||||
(Radersma, 2022) | ✓ | sustainability guidelines | ER | BB | 3 | ✓ | generic sw | ✓ | ✓ | ✓ | CP | 2018 | |||||||||||||
(Raisian et al., 2021) | ✓ | sustainability guidelines | GP | BB | 2 | ✓ | HPC | ✓ | JA | 2022 | |||||||||||||||
(Sharma et al., 2015) | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | CP | 2021 | ||||||||||||||
(Shenoy and Eeratta, 2011) | ✓ | ✓ | sustainability guidelines | GP | BB | 3 | ✓ | ✓ | generic sw | ✓ | JA | 2022 | |||||||||||||
(Siegmund et al., 2022) | ✓ | ✓ | measurement solutions | AT | BB | 4 | ✓ | ✓ | ML | ✓ | ✓ | ✓ | CP | 2023 | |||||||||||
(Sikand et al., 2023) | ✓ | sustainability guidelines | GP | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2011 | ||||||||||
(Simon et al., 2023) | ✓ | measurement solutions | ET | BB | 4 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2023 | ||||||||||
(Tee et al., 2014a) | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2014 | |||||||||
(Tee et al., 2014b) | ✓ | ✓ | reference models | RM | BB | 2 | ✓ | ✓ | generic sw | ✓ | ✓ | ✓ | ✓ | ✓ | CP | 2014 | |||||||||
(Wedyan et al., 2023) | ✓ | sw improvement techniques | TE | WB | 2 | ✓ | generic sw | ✓ | CP | 2023 | |||||||||||||||
(Wei et al., 2023) | ✓ | sw improvement techniques | ML | BB | 4 | ✓ | ✓ | ML | ✓ | ✓ | CP | 2023 | |||||||||||||
(Zhang et al., 2023) | ✓ | sw improvement techniques | OP | BB | 4 | ✓ | ✓ | edge computing | ✓ | ✓ | ✓ | ✓ | ✓ | JA | 2023 |
@@ 3.1. Sustainability Guidelines
In this section, we focus on the portion of the surveyed corpus that proposes sustainability-related guidelines to design and develop sustainable software. We start by illustrating those studies that focus on reducing energy consumption of running software. We then move to carbon-aware proposals, and finally discuss those works that are both carbon- and energy-aware. We organise our discussion by target applications domains.
@@
Energy-awareness. Targeting generic software, Kharchenko and Gorbenko (Kharchenko and Gorbenko, 2013) provide a set of good practices in the form of procedural steps intended for sustainability engineers to implement green management of system components. Such guidelines account for different aspects of system components (e.g., threats, fault-tolerance) and enable identifying existing pathological and evolutionary chains among those aspects. From the analysis of such chains it is possible to determine how to minimise the system’s resource (and energy) consumption, while improving a set of customisable quality metrics (e.g., reliability, availability, security, resilience).
@@
On a different line, addressing generic software architects, Ponsard et al. (Ponsard et al., 2018) conduct an experience report of an industrial case study about building and putting in operation a shared pick-up service. They show how to conduct the requirement analysis phase by also considering sustainability goals and their intertwining with the functional requirements set by other stakeholders of a software system. By distinguishing between green by software and green in software requirements, they take implementation and architectural decisions at different levels (e.g., choice of efficient programming language and algorithms, usage of lightweight virtualisation) and define KPIs to monitor for improving future releases.
@@
Moving to approaches that consider HPC as their application domain, Radersma (Radersma, 2022) lists a set of good practices for software developers to implement more environmentally sustainable and maintainable software artefacts, e.g., by adopting green programming languages, avoiding unnecessary data accesses in favour or computing results on-the-fly, employing state-of-the-art HPC hardware, and relying on cluster power management software to optimise energy consumption. Considering instead an experience report on ML, Brownlee et al. (Brownlee et al., 2021) rely on exhaustive and genetic search to (sub-)optimally tune the hyper-parameters of multi-layer perceptron networks. From a software developer perspective, they investigate the relations among different configurations, accuracy and energy consumption both in the training and inference phases on a set of benchmark networks. As a result, in their benchmark, they reduce energy consumption by 30–50% with a decrease in accuracy below 1%.
@@
Carbon-awareness. Still in the ML domain, Pan and Rajan (Pan and Rajan, 2022) are the sole to discuss good practices for software architects to directly reduce carbon emissions produced by convolutional neural networks. Instead of training convolutional neural networks from scratch, they recommend exploiting model decomposition – having each module capable of classifying a single output class – together with transfer, one-shot, and few-shot learning to favour reusability and replaceability and (consequently) to reduce carbon emissions related to model (re)training.
@@
Carbon- and energy-awareness. We then survey proposals that consider reducing both energy and carbon footprints in the design and development of software. In collaboration with an industry partner and targeting software developers, Fernandez and Lago (Condori-Fernández and Lago, 2018) identify a set of domain-generic green strategies and link them to the prioritisation of software quality requirements (e.g. usability, maintainability). Their objective is threefold as they aim at raising people-awareness on software sustainability, making service operations environmentally sustainable, and support sustainable development processes, by emphasising the need to consider sustainability throughout the software design phase and beyond. On a similar line, (Shenoy and Eeratta, 2011) suggest that quality and sustainability should be considered equally important throughout the whole software life cycle. They propose a set of enhancement practices for software developers to drive green software development.
@@
Still targeting generic software, Mehra et al. (Mehra et al., 2023) report on their experience of using a static analysis tool to identify code inefficiencies and act upon them suitably. Their experience report, clearly intended for application developers, showcases the potential of static analysis in detecting energy-inefficient patterns, assessing their impact, and providing remediation strategies. They indeed successfully refactor a large application codebase from an industrial use case. Siegmund et al. (Siegmund et al., 2022) analyse the relationship between software configuration and energy consumption. They propose ML-based code analyses methods to identify energy and carbon inefficiencies in software, from the developer’s perspective. On the other hand, they enable software users to select energy-optimised configurations for their software according to a black-box model. Their model accounts for non-monotonic energy consumption in the configuration space, which enables relying non-linear functions to model the system’s behaviour.
@@
Building upond the work by Shenoy and Eeratta (Shenoy and Eeratta, 2011), Chauhan and Saxena (Chauhan and Saxena, 2013) propose the Green Cloud Framework (GCF) as a set of best practices to handle software development lifecycle in cloud computing settings. Their work accounts for recommendations from organisations like Greenpeace to use clean and renewable energy sources in software development to make software development energy-aware and carbon-aware. Specifically, GCF is intended for software providers, developers and end-users, to make Service Level Agreements (SLAs) include guarantees sustainability metrics (i.e. carbon emissions and energy consumption). Differently, (Shenoy and Eeratta, 2011) mainly focuses on software developers, without envisioning the key role of SLAs between application providers and end-users.
@@
Speaking of ML, Patterson et al. (Patterson et al., 2022) discuss the energy consumption and carbon footprint of machine learning training and inference, also considering embodied carbon emissions. They eventually propose the so-called 4M best practices for software developers, which involve choosing greener energy locations for ML workloads (Map), using newest hardware possible (Machine), relying on cloud resources to save on environmental footprint (Mechanisaton), and exploit state-of-the-art lightweight neural networks (Models). Similarly, Fu et al. (Fu et al., 2021) provide good practices for software developers to compute carbon emissions costs for the search phase of a computer vision architecture based on deep learning over its whole life cycle.
@@ 3.2. Reference models
In this section, we focus on the largest portion of our surveyed corpus, which proposes reference models for designing and developing carbon-aware applications. As in the previous section, we first report on contributions that consider energy as the main driver to reduce carbon emissions, then move to those focussing on direct carbon reduction and, finally, we discuss proposals that are both energy- and carbon-aware.
@@
Energy-awareness. Sharma et al. (Sharma et al., 2015) and Mohankumar and Kumar (Mohankumar and Kumar, 2016) introduce a reference model for software developers, by relying and extending the software life cycle model of (Kneuper, 2018). Sharma et al. (Sharma et al., 2015) consider an iterative software development life cycle made of six stages, i.e., requirement analysis, green analysis, design, implementation, testing, and maintenance. The green analysis phase cyclically connects to all others and acts as a checkpoint entity to monitor and continuously enhance the sustainability of all other phases according to set guidelines. Also following an iterative approach, (Mohankumar and Kumar, 2016) only consider a four-stage life cycle made of requirements analysis, design, coding and testing. For the requirements analysis, they propose to use questions, guidelines, and generic templates to match customers’ needs with sustainable technology choices, use well-timed metrics data for the purpose, and formal methods to avoid ambiguities. Based on requirements specification, developers should write the documentation concerning green and sustainable templates for designing all application components. Then, throughout the implementation phase they should always adopt best practices, e.g. avoid unnecessary classes, closing database connections after use, avoid stand-alone installation in favour of on-demand remote access.
@@
The reference model by Kern et al. (Kern et al., 2013) goes towards defining an eco-labelling to drive software developers in the design and development phases. It can be applied by analysing which tasks or functions of the software are most frequently used and then designs case studies accordingly, also considering servers’ energy consumption. This way of proceeding allows developers and end-users to compare and contrast different software implementation by determining the most sustainable one. Subsequent proposals by Kern et al. are also meant for IT operators (Kern et al., 2013) and for sustainability engineers and project owners (Kern et al., 2018).
@@
Carbon-awareness. To reduce carbon footprint and to raise green and sustainable awareness in generic software, Lami et al. (Lami et al., 2014, 2013) define a methodological approach for product owners. They use such an approach to derive sustainability indicators for the whole life cycle of a software project and use them to populate bi-dimensional matrices, known as hierarchical software templates. Such templates are used to explore three levels of environmentally effects, viz. direct, indirect, and systemic, also associating direct effects with their measurements and success factors. Last, in compliance with ISO/IEC 33000, Lami et al. (Lami et al., 2014) define the assessment of the sustainability software process via a framework made of six process sustainability levels, viz. incomplete, performed, managed, defined, measured, and optimised.
@@
Energy-awareness and carbon-awareness. All studies on reference models considering both energy and carbon reduction are intended for the generic software domain with a black-box approach. For example, Tee et al. (Tee et al., 2014a, b) define a reference model for both software developers and IT operators, by proposing a procedural method for making software development life cycle more sustainable. They aim at gathering, assembling and storing of the information for the knowledge management of a green and sustainable software development life cycle. They collect information on existing findings, processes and activities, and techniques in software development, and propose a knowledge management model and step-by-step guidelines to make software sustainable. They then map techniques to actionable tasks to minimise power consumption and carbon footprint while developing software. Following up on the above, Tee et al. (Tee et al., 2014a) extend the knowledge-management model into a green-management model to be shared among all stakeholders to build an action-oriented and collaborative community of practice.
@@
Jayanthi et al. (Jayanthi et al., 2021) and Dick et al. (Dick et al., 2013) provide a reference model for software developers, sustainability engineers, and end-users by defining green and sustainable software development life cycle steps integrated within Agile operations. In particular, Jayanthi et al. (Jayanthi et al., 2021) advocate organisations should introduce a sustainability department including quality assurance engineers for the requirements analysis phase, user experience designers for the design phase, certified green computing developers for the coding and testing phase, and managers for the overall success of the sustainable software deployment. Conversely, Dick et al. (Dick et al., 2013) add a specific sustainable development stage to the software life cycle model. They associate SCRUM activities to their driving role among sustainability engineers and software developers, and embed sustainability reporting and retrospective in the Agile development model.
@@
On a similar line, extending their previous effort (Lami et al., 2014) to consider both energy and carbon footprints, Lami and Buglione (Lami and Buglione, 2012) provide a reference model for project owners to enact measurements by considering a cause-effect diagram analysis on the infrastructure, people, software process and product. The model has to create a suitable measurement plan which integrates all needed information elements for specifically addressing the sustainability improvement throughout software development. They add sustainability in project management as a first-class citizen like quality, time, and budget. With an akin objective, Ibrahim et al. (Ibrahim et al., 2021; Ibrahim et al., 2022) provide a reference model for both software developers and end-users. They focus holistically on software product development, as well as the related hardware disposal and recycling. Their approach also relies on protocols for (semi-)structured interviews with involved stakeholders to collect data useful for process mining from a sustainability viewpoint. The proposed model aims at achieving resource and energy efficiency to pursue long-term improvement on software sustainability, with a particular focus on reducing waste (e.g., over-requirements, miscommunication, rework and lack of knowledge). Ibrahim et al. (Ibrahim et al., 2021) also showcases the proposed approach over real industrial use cases.

### v2

@@ 2.5. Replication Package
To enable repeating our study, as well as verifying our findings, we publicly released a replication package on Zenodo (Danushi et al., 2025). The replication package includes the intermediate artifacts produced during the search and selection of studies to be considered, as well as their final classification. The selected studies can instead be accessed by relying on the bibliographic information provided by the references available in Table 2.
@@ 3. Overviewing the State-of-the-Art Solutions
[…]
@@ 3.1. Sustainability Guidelines
In this section, we focus on the portion of the surveyed corpus that proposes sustainability-related guidelines to design and develop sustainable software. We start by illustrating those studies that focus on indirectly reducing the carbon footprint of running software. We then move to those directly considering the carbon footprint of software, and finally discuss those works that follow a hybrid approach. We organise our discussion by target application domain.
@@
Indirect. For generic software, Kharchenko and Gorbenko (Kharchenko and Gorbenko, 2013) provide a set of good practices in the form of procedural steps intended for sustainability engineers to implement green management of system components. Such guidelines account for different aspects of system components (e.g., threats, fault-tolerance) and enable identifying existing pathological and evolutionary chains among those aspects. From the analysis of such chains it is possible to determine how to minimise the system’s resource (and energy) consumption, while improving a set of customisable quality metrics (e.g., reliability, availability, security, resilience).
@@
On a different line, addressing generic software architects, Ponsard et al. (Ponsard et al., 2018) conduct an experience report of an industrial case study about building and putting in operation a shared pick-up service. They show how to conduct the requirement analysis phase by also considering sustainability goals and their intertwining with the functional requirements set by other stakeholders of a software system. By distinguishing between green by software and green in software requirements, they take implementation and architectural decisions at different levels (e.g., choice of efficient programming language and algorithms, use of lightweight virtualisation) and define key performance indicators to monitor for improving future releases.
@@
Moving to approaches that consider HPC as their application domain, Radersma (Radersma, 2022) lists a set of good practices for software developers to implement more environmentally sustainable and maintainable software artefacts, e.g., by adopting green programming languages, avoiding unnecessary data accesses in favour or computing results on-the-fly, employing state-of-the-art HPC hardware, and relying on cluster power management software to optimise energy consumption. Considering instead an experience report on ML, Brownlee et al. (Brownlee et al., 2021) rely on exhaustive and genetic search to (sub-)optimally tune the hyperparameters of multi-layer perceptron networks. From a software developer’s perspective, they investigate the relations among different configurations, accuracy and energy consumption in both the training and inference phases on a set of benchmark networks. As a result, in their benchmark, they reduce energy consumption by 30–50% with a decrease in accuracy below 1%.
@@
Direct. Still in the ML domain, Pan and Rajan (Pan and Rajan, 2022) are the sole to discuss good practices for software architects to directly reduce carbon emissions produced by convolutional neural networks. Instead of training convolutional neural networks from scratch, they recommend exploiting model decomposition – having each module capable of classifying a single output class – together with transfer, one-shot, and few-shot learning to favour reusability and replaceability and (consequently) reduce the carbon emissions associated with model (re)training.
@@
Hybrid. In collaboration with an industry partner and targeting software developers, Fernandez and Lago (Condori-Fernández and Lago, 2018) identify a set of domain generic green strategies and link them to the prioritisation of software quality requirements (e.g. usability, maintainability). Their aim is threefold: to raise people’s awareness of software sustainability, to make service operations environmentally sustainable, and to support sustainable development processes, by emphasising the need to consider sustainability throughout the software design phase and beyond. Similarly, Shenoy and Eeratta (Shenoy and Eeratta, 2011) suggest that quality and sustainability should be considered equally important throughout the whole software lifecycle. They propose a set of improvement practices for software developers to drive green software development.
@@
Still targeting generic software, Mehra et al. (Mehra et al., 2023) report on their experience of using a static analysis tool to identify code inefficiencies and act upon them suitably. Their experience report, clearly intended for application developers, showcases the potential of static analysis in detecting energy-inefficient patterns, assessing their impact, and providing remediation strategies. They indeed successfully refactor a large application codebase from an industrial use case. Siegmund et al. (Siegmund et al., 2022) analyse the relationship between software configuration and energy consumption. They propose ML-based code analysis methods to identify energy and carbon inefficiencies in software, from the developer’s perspective. On the other hand, they enable software users to select energy-optimised configurations for their software according to a black-box model. Their model accounts for non-monotonic energy consumption in the configuration space, which enables relying non-linear functions to model the system’s behaviour.
@@
Building upon the work by Shenoy and Eeratta (Shenoy and Eeratta, 2011), Chauhan and Saxena (Chauhan and Saxena, 2013) propose the Green Cloud Framework (GCF) as a set of best practices to handle software development lifecycle in cloud computing settings. Their work accounts for recommendations from organisations like Greenpeace to use clean and renewable energy sources in software development to make software development energy-aware and carbon-aware. Specifically, GCF is intended for software providers, developers and end-users, to make Service Level Agreements (SLAs) include guarantees of sustainability metrics (i.e. carbon emissions and energy consumption). Differently, (Shenoy and Eeratta, 2011) mainly focuses on software developers, without envisioning the key role of SLAs between application providers and end-users.
@@
Speaking of ML, Patterson et al. (Patterson et al., 2022) discuss the energy consumption and carbon footprint of machine learning training and inference, also considering embodied carbon emissions. They eventually propose the so-called 4M best practices for software developers, which involve choosing greener energy locations for ML workloads (Map), using the latest possible hardware (Machine), relying on cloud resources to reduce the environmental footprint (Mechanisaton), and exploit state of the art lightweight neural networks (Models). Similarly, Fu et al. (Fu et al., 2021) provide good practices for software developers to compute carbon emissions costs for the search phase of a computer vision architecture based on deep learning over its entire lifecycle.
@@ 3.2. Reference models
In this section, we focus on the largest part of our surveyed corpus, which proposes reference models for the design and development of carbon-aware applications. As in the previous section, we first report on contributions that indirectly reduce carbon emissions, and then move to those that focus on direct carbon reduction. Finally, we discuss hybrid proposals.
@@
Indirect. Sharma et al. (Sharma et al., 2015) and Mohankumar and Kumar (Mohankumar and Kumar, 2016) introduce a reference model for software developers, by relying and extending the software lifecycle model of (Kneuper, 2018). Sharma et al. (Sharma et al., 2015) consider an iterative software development lifecycle made of six stages, i.e., requirement analysis, green analysis, design, implementation, testing, and maintenance. The green analysis phase cyclically connects to all others and acts as a checkpoint entity to monitor and continuously enhance the sustainability of all other phases according to set guidelines. Also following an iterative approach, Mohankumar and Kumar (Mohankumar and Kumar, 2016) only consider a four-stage lifecycle consisting of requirements analysis, design, coding and testing. For the requirements analysis, they propose to use questions, guidelines, and generic templates to match customers’ needs with sustainable technology choices, use well-timed metrics data for the purpose, and formal methods to avoid ambiguities. Based on requirements specification, developers should write the documentation concerning green and sustainable templates for designing all application components. Then, throughout the implementation phase they should always adopt best practices, e.g. avoid unnecessary classes, closing database connections after use, avoid stand-alone installation in favour of on-demand remote access.
@@
The reference model by Kern et al. (Kern et al., 2013) goes towards defining an eco-labelling to drive software developers in the design and development phases. It can be applied by analysing which tasks or functions of the software are most frequently used and then designing case studies accordingly, also considering servers’ energy consumption. This way of proceeding allows developers and end-users to compare and contrast different software implementations by determining the most sustainable one. Subsequent proposals by Kern et al. are also aimed at IT operators (Kern et al., 2013) and for sustainability engineers and project owners (Kern et al., 2018).
@@
Direct. To reduce carbon footprint and to raise green and sustainable awareness in generic software, Lami et al. (Lami et al., 2014, 2013) define a methodological approach for product owners. They use such an approach to derive sustainability indicators for the whole lifecycle of a software project and use them to populate bi-dimensional matrices, known as hierarchical software templates. Such templates are used to explore three levels of environmental effects, viz. direct, indirect, and systemic, also associating direct effects with their measurements and success factors. Last, in compliance with ISO/IEC 33000, Lami et al. (Lami et al., 2014) define the assessment of six process sustainability levels, viz. incomplete, performed, managed, defined, measured, and optimised.
@@
Hybrid. All studies on hybrid reference models are intended for the generic software domain with a black-box approach. For example, Tee et al. (Tee et al., 2014a, b) define a reference model for both software developers and IT operators, by proposing a procedural method for making software development lifecycle more sustainable. They aim at gathering, assembling and storing of the information for the knowledge management of a green and sustainable software development lifecycle. They collect information on existing findings, processes and activities, and techniques in software development, and propose a knowledge management model and step-by-step guidelines to make software sustainable. They then map techniques to actionable tasks to minimise power consumption and carbon footprint while developing software. Following up on the above, Tee et al. (Tee et al., 2014a) extend the knowledge-management model into a green-management model to be shared among all stakeholders to build an action-oriented and collaborative community of practice.
@@
Jayanthi et al. (Jayanthi et al., 2021) and Dick et al. (Dick et al., 2013) provide a reference model for software developers, sustainability engineers, and end-users by defining green and sustainable software development lifecycle steps integrated within Agile operations. In particular, Jayanthi et al. (Jayanthi et al., 2021) argue that organisations should establish a sustainability department that includes quality assurance engineers for the requirements analysis phase, user experience designers for the design phase, certified green computing developers for the coding and testing phase, and managers for the overall success of the sustainable software deployment. Conversely, Dick et al. (Dick et al., 2013) add a specific sustainable development stage to the software lifecycle model. They associate SCRUM activities to their driving role among sustainability engineers and software developers, and embed sustainability reporting and retrospective in the Agile development model.
@@
On a similar line, extending their previous effort (Lami et al., 2014), Lami and Buglione (Lami and Buglione, 2012) provide a reference model for project owners to enact measurements by considering a cause-effect diagram analysis on the infrastructure, people, software process and product. The model has to create a suitable measurement plan that integrates all needed information elements to specifically address the sustainability improvement throughout software development. They add sustainability to project management as a first-class citizen like quality, time, and budget. With an akin objective, Ibrahim et al. (Ibrahim et al., 2021, 2022) provide a reference model for both software developers and end-users. They focus holistically on software product development, as well as the related hardware disposal and recycling. Their approach also relies on protocols for (semi-)structured interviews with involved stakeholders to collect data useful for process mining from a sustainability viewpoint. The proposed model aims at achieving resource and energy efficiency to pursue long-term improvement of software sustainability, with a particular focus on reducing waste (e.g., over-requirements, miscommunication, rework, and lack of knowledge). Ibrahim et al. (Ibrahim et al., 2021) also showcase the proposed approach over real industrial use cases.